A business analyst wants to explore the relationship between the sales for a product (in thousands of dollars) and the possible factors; price (in dollars) and the promotional expenditure(in hundred of dollars). Results of the partial multiple regression analysis are presented below.

Regression Analysis: Sales versus Price, Promotion

The regression equation is

Sales = 12.9 - 6.50 Price + 0.0207 Promotion

Predictor     Coef SE Coef       T      P   

Constant   12.8638   0.2425   53.05 0.000

Price -6.4979   0.3575 -18.18 0.000

Promotion 0.02069 0.03811    0.54 0.611

Analysis of Variance

Source          DF      SS MS F P

Regression             9.1578 4.5789 0.000

Residual Error 0.1222 0.0244

Total 7 9.2800

What is the F test statistic value for the regression?

Select one:

a. 156.75

b. 21.34

c. 187.66

d. 98.56

A business analyst wants to explore the relationship between the sales for a product (in thousands of dollars) and the possible factors; price (in dollars) and the promotional expenditure(in hundred of dollars). Results of the partial multiple regression analysis are presented below.

Regression Analysis: Sales versus Price, Promotion

The regression equation is

Sales = 12.9 - 6.50 Price + 0.0207 Promotion

Predictor     Coef SE Coef       T      P   

Constant   12.8638   0.2425   53.05 0.000

Price -6.4979   0.3575 -18.18 0.000

Promotion 0.02069 0.03811    0.54 0.611

Analysis of Variance

Source          DF      SS MS F P

Regression             9.1578 4.5789 0.000

Residual Error 0.1222 0.0244

Total 7 9.2800

What is the predicted sales (estimated average sales) if the price charged is 60 cents and the promotional expenditure is 300 dollars?

Select one:

a. $9,062

b. $90,621

c. $9.062

d. $906

In multiple regression, the ________ procedure permits variables to enter and leave the model at different stages of its development.

Select one:

a. stepwise regression

b. residual analysis

c. backward elimination

d. Best-Subsets Approach
Answer #1

1)F test statistic value for the regression = 4.5789/0.0244 =187.66

2)

predicted sales =(12.9-(6.5*0.6)+0.0207*3)*1000 =9062

3)

In multiple regression, the stepwise regression procedure permits variables to enter and leave the model at different stages of its development.